40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 11 11:48:43 2020, gautam, Update, VAC, Vac failure vacFailure.png
    Reply  Thu Jun 11 16:14:03 2020, gautam, Update, VAC, Vac failure - probable cause is serial comm glitch TP2_tempGlitch.pngPSL_shutterClosed.pngTP2tempGlitches.pdf
       Reply  Thu Jun 11 17:35:34 2020, gautam, Update, VAC, Pumpspool UPS needs battery replacement 
          Reply  Fri Jun 12 01:23:32 2020, Koji, Update, VAC, Pumpspool UPS needs battery replacement 
             Reply  Fri Jun 12 17:32:40 2020, Koji, Update, VAC, Pumpspool UPS needs battery replacement Screen_Shot_2020-06-12_at_17.22.23.png
                Reply  Fri Jun 12 19:02:52 2020, gautam, Update, VAC, Pumpspool UPS needs battery replacement 
                   Reply  Fri Jun 12 19:23:56 2020, Koji, Update, VAC, Pumpspool UPS needs battery replacement 
                      Reply  Fri Jun 12 19:33:31 2020, gautam, Update, VAC, Pumpspool UPS needs battery replacement 
                         Reply  Tue Jun 16 13:35:03 2020, Jon, Update, VAC, Temporary vac fix / IFO usable again Pumpdown.png
                            Reply  Wed Jun 17 16:27:51 2020, gautam, Update, VAC, Questions/comments on vacuum 
                               Reply  Thu Jun 18 11:00:24 2020, Jon, Update, VAC, Questions/comments on vacuum 
                                  Reply  Thu Jun 18 12:00:36 2020, gautam, Update, VAC, Questions/comments on vacuum 
                                     Reply  Thu Jun 18 14:13:03 2020, Jon, Update, VAC, Questions/comments on vacuum 
                                        Reply  Thu Jun 18 15:46:34 2020, gautam, Update, VAC, Questions/comments on vacuum 
                                           Reply  Fri Jun 19 07:40:49 2020, Jon, Update, VAC, Questions/comments on vacuum 
                                              Reply  Fri Jun 19 09:57:35 2020, gautam, Update, VAC, Questions/comments on vacuum 
Message ID: 15396     Entry time: Fri Jun 12 17:32:40 2020     In reply to: 15394     Reply to this: 15397
Author: Koji 
Type: Update 
Category: VAC 
Subject: Pumpspool UPS needs battery replacement 

Jon and Koji remotely supported Jordan's resetting the TP2 controller.

Here is the instruction by Jon
From the operator's console in front of the vac rack:
  1. Open a terminal window (click the LXTerminal icon on the desktop)
  2. Type "control" + enter to open the vac controls screen
  3. Toggle all the open valves closed (edit by KA: and manually close RV2 by rotating the gate valve handle )
  4. Turn OFF TP2 by clicking the "Off' button. Make sure the status changes and the rotation speed falls to zero (you'll also hear the pump spinning down) 
  5. The other pumps (TP1, TP3) can be left running
  6. Once TP2 has stopped spinning, go to the back of the rack and locate the ethernet cable running from the back of the TP2 controller to the IOLAN server (near the top of the rack). Disconnect and reconnect the cable at each end, verifying it is firmly locked in place.
  7. From the front of the rack, power down the TP2 controller (I don't quite remember for the Agilent, but you might have to move the slider on the front from "Remote" to "Local" first)
  8. Wait about 30 seconds, then power it back on. If you had to move the slider to shut it down, revert it back to the "Remote" position.
  9. Go back to the controls screen on the console. If the pump came back up and is communicating serially again, its status will say something other than "NO COMM"
  10. Turn TP2 back on. Verify that it spins up to its nominal speed (66 kRPM)
  11. At this point you can reopen any valves you initially closed (any that were already closed before, leave closed)

TP2 was stopped and at this moment the glitches were gone. Jordan powercycled the TP2 controller and we brought up the TP2 back at the full speed.
However, the glitches came back as before. Obviously we can't go on from here, and we've decided to stop the recovery process here today.


- We left TP1/2/3 running while the valves including RV2 were closed.

- When Jordan is back in the lab next week, we'll try to use TP3 as the backing of TP1 so that we can resume the main volume pumping.

- Currently, TP3 does not have interlocking and that is a risk. Jon is going to implement it.

- Meanwhile, we will try to replace the controller of TP2. We are supposed to have this in the lab. Ask Chub about the location.

- Once we confirm the stability of the diagnostic signals for TP2, we will come back to the nominal pumping scheme.

Attachment 1: Screen_Shot_2020-06-12_at_17.22.23.png  115 kB  Uploaded Fri Jun 12 18:40:08 2020  | Hide | Hide all
Screen_Shot_2020-06-12_at_17.22.23.png
ELOG V3.1.3-